OpenStudy (anonymous):

Are the two triangles below similar? Yes; they have congruent corresponding angles Yes; they have proportional corresponding sides No; they do not have proportional corresponding sides No; they do not have congruent corresponding angles

OpenStudy (anonymous):

Yes; they have congruent corresponding angles Yes; they have proportional corresponding sides No; they do not have proportional corresponding sides No; they do not have congruent corresponding angles

OpenStudy (anonymous):

I think its No; they do not have proportional corresponding sides @MOj0jojo

OpenStudy (anonymous):

I'm trying to solve for all the sides, one sec.

OpenStudy (cwrw238):

the last option is correct

OpenStudy (anonymous):

ok what do you think @cwrw238

OpenStudy (cwrw238):

if you work out the value of the third angle in the smaller triangle you'll see that thecorresponding angles are not equal

OpenStudy (anonymous):

Its either of the last two options.

OpenStudy (anonymous):

not sure which one is correct though.

OpenStudy (anonymous):

T and W are both the same degree and yea Ik its between the last 2 I just don't know which one it is.

OpenStudy (cwrw238):

3rd angle in small triangle is 180 - (70 = 280 = 180 - 98 = 82

OpenStudy (anonymous):



OpenStudy (cwrw238):

* that was (70 + 28)

OpenStudy (cwrw238):

right - from the diagram you see its corresponding angles which are not congruent

OpenStudy (anonymous):

so its no; they do not have congruent corresponding angles

OpenStudy (cwrw238):

yes last option

OpenStudy (anonymous):

thank you guys so much would you guys mind helping me with others or checking it because im not good in geometry

OpenStudy (cwrw238):

by definition similar triangles have congruent corresponding angles